# Factor Augmented Tensor-on-Tensor Neural Networks

### v1

@@ A.5 Tucker Decomposition and iterative TIPUP (iTIPUP) Algorithm
A.6 Proof of Proposition 3.1
Let $L_{k}^{(m)}$ be the loss for $\hat{A}_{k}^{(m)}$ at the $m$-th iteration,
We outline the proof as follows.
(i) Initialization.
Define
The “noiseles” version of TIPUP1 is
which is of rank $r_{1}$. Then, by Davis-Kahan sin $\Theta$ theorem (Lemma A.2),
To bound the numerator on the right hand side of (S.13), we write
${\rm TIPUP}_{1}-\Theta_{1}=\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}{\cal E}_{t}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}A_{2}{\cal F}_{t}^{\top}A_{1}^{\top}+\frac{1}{n}\sum_{t=1}^{n}{\cal E}_{t}{\cal E}_{t}^{\top}:=\Delta_{1}+\Delta_{2}+\Delta_{3}.$
For $\Delta_{3}$, by Lemma A.7,
$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}\mathcal{E}_{t}^{\top}\right\|_{2}\geq d_{1}+2\sqrt{d_{1}d_{2}n}+2x(\sqrt{d_{1}}+\sqrt{d_{2}n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
By setting $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
Thus for $u_{i}$ and $v_{i}$ with $\|u_{i}\|_{2}=\|v_{i}\|_{2}=1$, $i=1,2$,
$n\|\Theta_{1}\|_{2}^{-1}\overline{\mathbb{E}}\left((u_{1}^{\top}\Delta_{1}v_{1}-u_{2}^{\top}\Delta_{1}v_{2})^{2}\right)$ $\leq\left(\|u_{1}-u_{2}\|_{2}\|v_{1}\|_{2}+\|u_{2}\|_{2}\|v_{1}-v_{2}\|_{2}\right)^{2}$
$\leq 2\left(\|u_{1}-u_{2}\|_{2}^{2}+\|v_{1}-v_{2}\|_{2}^{2}\right)$
where $\xi$ and $\zeta$ are i.i.d $N(0,I_{d_{1}})$ vectors. As $\Delta_{1}$ is a $d_{1}\times d_{1}$ Gaussian matrix under $\overline{\mathbb{E}}$, the Sudakov-Fernique inequality (Lemma A.5) yields
$\sqrt{n}\|\Theta_{1}\|_{2}^{-1/2}\overline{\mathbb{E}}\|\Delta_{1}\|_{2}\leq\sqrt{2}\mathbb{E}\sup_{\|u\|_{2}=\|v\|_{2}=1}(u^{\top}\xi+v^{\top}\zeta)=\sqrt{2}\mathbb{E}\left(\|\xi\|_{2}+\|\zeta\|_{2}\right)\leq 2\sqrt{2d_{1}}.$
It follows that
Elementary calculation shows that $\|\Delta_{1}\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}^{1/2}$-Lipschitz function. Then, by Gaussian concentration inequality for Lipschitz functions (Lemma A.6),
With $x\asymp\sqrt{d_{1}}$, with probability at least $1-e^{-d_{1}}$,
(ii) Iterative refinement.
[…]
$\widehat{A}_{1}^{(m+1)}=\text{LSVD}_{r_{1}}(\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})),\quad\widehat{P}_{1}^{(m+1)}=\widehat{A}_{1}^{(m+1)}\widehat{A}_{1}^{(m+1)^{\top}}.$ (S.15)
The “noiseless” version of this update is given by
giving error-free “estimates”,
$\widehat{A}_{1}^{(m+1)}=\text{LSVD}_{r_{1}}(\Theta_{1}(\widehat{A}_{2}^{(m)})),$
where $\Theta_{1}(\widehat{A}_{2}^{(m)})$ is of rank $r_{1}$. Thus, by Davis-Kahan sin $\Theta$ theorem (Lemma A.2)
To bound the numerator on the right-hand side of (S.16), we write
where for any $M_{2}\in\mathbb{R}^{d_{2}\times d_{2}}$,
$\|\text{TIPUP}_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(\widehat{A}_{2}^{(m)})\|_{2}\leq\|\text{TIPUP}_{1}(A_{2})-\Theta_{1}(A_{2})\|_{2}+L^{(m)}(2K-2)\sum_{i=1}^{3}\|\Delta_{i}(M_{2})\|_{\rm S},$
where
We will also prove by induction, the denominator in (S.16) satisfies
Define the ideal version of the ratio in (S.16) as
$L_{k}^{(\text{ideal})}=\frac{\|\text{TIPUP}_{1}(\odot_{j\neq k}A_{j})-\Theta_{k}(\odot_{j\neq k}A_{j})\|_{2}}{\sigma_{r_{1}}[\Theta_{k}(\odot_{j\neq k}A_{j})]},\quad L^{(\text{ideal})}=\max_{1\leq k\leq K}L_{k}^{(\text{ideal})}.$
Putting together (S.13),(S.14),(S.16),(S.17),(S.18) and (S.19), we have
$L_{k}^{(m+1)}\leq L_{k}^{(\text{ideal})}+L^{(m)}\rho,$
with $\rho<1$. By induction,
$L^{(m+1)}\leq(1+\rho+\dots+\rho^{m})L^{(m)}+\rho^{m+1}L^{(0)}.$
We achieve the desired results.
Step 1.
We prove (S.17) for the $\|\Delta_{1}(M_{2})\|_{\rm S}$.
By Lemma A.3 (ii), there exists $\widetilde{M}^{(\ell,\ell^{\prime})}\in\mathbb{R}^{d_{2}\times d_{2}}$ of the forms $W_{\ell}Q_{\ell^{\prime}}^{\top}$ with $W_{\ell}\in\mathbb{R}^{d_{2}\times r_{2}}$, $Q_{\ell^{\prime}}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell,\ell^{\prime}\leq N_{d_{2}r_{2},1/8}=17^{d_{2}r_{2}}$, such that $\|\widetilde{M}^{(\ell,\ell^{\prime})}\|_{2}\leq 1$, $\text{rank}(\widetilde{M}^{(\ell,\ell^{\prime})})\leq r_{2}$ and
Elementary calculation shows that $\|\Delta_{1}(\widetilde{M}^{(\ell,\ell^{\prime})})\|_{2}$ is a $\sqrt{n}\|\Theta_{1}\|_{2}$-Lipschitz function of $\mathcal{E}_{1},\ldots,\mathcal{E}_{n}$. Employing similar argument as the initialization (Sudakov-Fernique inequality, Lemma A.5), we have
Then, by Gaussian concentration inequality for Lipschitz functions (Lemma A.6),
¯$\overline{\mathbb{P}}\left(\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}\widetilde{M}^{(\ell,\ell^{\prime})}\mathcal{E}_{t}^{\top}\right\|_{2}-\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}\geq\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq e^{-x^{2}/2}.$
Hence,
¯$\overline{\mathbb{P}}\left(\|\Delta_{1}(M_{2})\|_{\rm S}\geq\sqrt{\frac{8d_{1}}{n}}\|\Theta_{1}\|_{2}^{1/2}+\frac{1}{\sqrt{n}}\|\Theta_{1}\|_{2}^{1/2}x\right)\leq N_{d_{2}r_{2},1/8}^{2}e^{-x^{2}/2}.$
Step 2.
Step 3.
Here we prove (S.18) for $\|\Delta_{3}(M_{2})\|_{\rm S}$. By Lemma A.3(iii), we can find $U_{2}^{(\ell)}\in\mathbb{R}^{d_{2}\times r_{2}}$, $1\leq\ell\leq N_{d_{2}r_{2},1/8}$ such that $\|U_{2}^{(\ell)}\|_{2}\leq 1$ and
By Lemma A.7,
$\mathbb{P}\left(\left\|\sum_{t=1}^{n}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}-n\mathbb{E}\mathcal{E}_{t}U_{2}^{(\ell)}U_{2}^{(\ell)\top}\mathcal{E}_{t}^{\top}\right\|_{2}\geq d_{1}+2\sqrt{d_{1}r_{2}n}+2x(\sqrt{d_{1}}+\sqrt{r_{2}n})+x^{2}\right)\leq 2e^{-x^{2}/2}.$
$\|\Delta_{3}(M_{2})\|_{\rm S}\lesssim\frac{d_{1}}{n}+\sqrt{\frac{d_{1}r_{2}}{n}}+\frac{d_{2}r_{2}}{n}+\sqrt{\frac{d_{2}r_{2}^{2}}{n^{2}}}\lesssim\frac{d_{1}+d_{2}}{n}+\sqrt{\frac{d_{1}+d_{2}}{n}}.$
Step 4.
Next, we prove (S.19). Note that
$\|\Theta_{1}(\widehat{A}_{2}^{(m)})-\Theta_{1}(A_{2})\|_{2}$ $\leq\left\|\frac{1}{n}\sum_{t=1}^{n}A_{1}F_{t}A_{2}^{\top}(\widehat{A}_{2}^{(m)}\widehat{A}_{2}^{(m)\top}-A_{2}A_{2}^{\top})A_{2}F_{t}^{\top}A_{1}^{\top}\right\|_{2}$
$\leq L^{(m)}\|\Theta_{1}\|_{2}.$
Thus, by Weyl’s inequality (Lemma A.1),
when $\min_{1\leq k\leq K}\frac{\sigma_{r_{k}}[\Theta_{k}(A_{2})]}{\|\Theta_{k}(A_{2})\|_{2}}\geq 2L^{(m)}$. We prove this condition by induction.
Appendix B Supplemental Results to Numerical Studies in Section 4
B.1 Supplemental Notation used in Simulation Studies
(1) For vectors $a_{1},\ldots,a_{K}$ of length $r_{1},\ldots,r_{K}$, the outer product
(2) Let $M_{1}\in\mathbb{R}^{r_{1}\times R},\ldots M_{K}\in\mathbb{R}^{r_{K}\times R}$ be $K$ matrices with the same column dimension (denoted by $R$), the notation $\llbracket M_{1},\cdots,M_{K}\rrbracket$ represents
where $m_{kr}$ is the $r$th column of $M_{k}$, $k=1,\ldots,K$, $r=1,\ldots,R$.
(3) For two tensor $\mathcal{A}:d_{1}\times\cdots\times d_{K}\times r_{1}\times\cdots\times r_{K}$ and $\mathcal{B}:r_{1}\times\cdots\times r_{K}\times p_{1}\times\cdots\times p_{q}$, the notation $\langle{\cal A},{\cal B}\rangle_{L}$ denote the contracted tensor product (Lock,, 2018)
$\langle\mathcal{A},\mathcal{B}\rangle_{L}[i_{1},\ldots,i_{K},j_{1},\ldots,j_{d}]=\sum_{l_{1}=1}^{r_{1}}\cdots\sum_{l_{K}=1}^{r_{K}}\mathcal{A}[i_{1},\dots,i_{K},l_{1},\dots,l_{K}]\mathcal{B}[l_{1},\dots,l_{K},j_{1},\dots,j_{d}].$ (S.22)
B.2 Supplemental Results for FAO Data Analysis
  • (i)
  • (ii)
Table S.2 and Table S.3 supplement Figure 3 by providing the numerical values of ground truth and predicted Yield metrics for the plot.
@@ B.2 Supplemental Results for FAO Data Analysis
Table S.2: Numerical values of ground-truth and predicted Yield of cirtus fruit in 26 European countries using livestock data of the same 26 countries - a supplement to Figure 3
@@
Note: Values are reported on the log-transformed scale.
Table S.3: Numerical values of ground-truth and predicted Area-harvested of cirtus fruit in 26 European countries using livestock data of the same 26 countries - a supplement to Figure 3
Country True Value FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
Austria 10.01 10.74 13.54
Bulgaria 9.36 11.81 13.65
Croatia 9.18 10.03 0.33
Cyprus 8.39 12.07 11.58
Czechia 10.07 9.54 7.65
Denmark 8.55 7.98 1.90
Estonia 9.99 10.53 13.61
Finland 12.14 14.00 11.41
France 12.44 11.48 9.40
Germany 10.07 11.60 10.08
Greece 9.65 12.03 12.86
Hungary 9.14 11.37 8.96
Ireland 10.84 12.24 9.91
Italy 11.82 11.87 9.86
Latvia 10.71 10.94 5.96
Lithuania 9.75 8.53 4.28
Luxembourg 10.05 8.26 3.16
Malta 6.42 -1.43 -0.02
Netherlands 9.65 10.44 7.39
Poland 11.98 12.27 15.52
Portugal 12.61 12.11 22.19
Romania 12.02 14.08 18.88
Slovakia 9.00 6.94 -0.97
Slovenia 8.20 6.61 0.83
Spain 11.17 7.22 14.15
Sweden 10.10 7.71 13.22
Note: Values are reported on the log-transformed scale.
B.3 Supplemental Information for NYC Taxi Data Analysis

### v2

@@ Appendix C Appendix C: An Extended Discussion
Figure S.1: Comparisons between ground-truth values and predicted Production of Treenuts (top row) and Yield of Oilcrops (bottom row) in South America using the crop data of 33 countries in East Asia, North America, and Europe. Values are plotted on the log-transformed scale. From left to right: Ground truth, FATTNN, TCN, LSTM, TRL, and Conv-TT-LSTM.
Figure S.2: Comparisons between ground-truth values and predicted Area-harvested (top row) and Yield (bottom row) of citrus fruit in 26 selected countries in Europe using the livestock data of the same 26 countries. Values are plotted on the log-transformed scale. From left to right: Ground truth, FATTNN, TCN, LSTM, TRL, and Conv-TT-LSTM.
Table S.2: Numerical values of ground-truth and predicted (a) Production of Treenuts and (b) yield Oilcrops in 13 South America countries using crop data of 33 countries in East Asia, North America, and Europe - a supplement to Figure 2.
(a) Predicted Production of Treenuts
Country Ground Truth FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
Argentina 11.77 14.77 4.30
Bolivia 10.84 11.15 6.48
Brazil 13.55 13.73 9.56
Chile 9.91 8.88 7.81
Colombia 11.35 10.66 6.51
Ecuador 10.51 11.32 7.67
French Guiana -9.21 -0.71 2.80
Guyana 7.30 4.34 3.98
Paraguay 9.32 9.68 6.11
Peru 11.20 10.40 7.46
Suriname 7.42 7.03 6.02
Uruguay 9.61 5.87 5.78
Venezuela 10.69 15.72 11.30
(b) Predicted Yield of Oilcrops
Country Ground Truth FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
Argentina 9.69 13.53 4.31
Bolivia 9.91 9.49 7.11
Brazil 7.99 9.91 6.48
Chile 10.18 10.30 8.30
Colombia 10.31 9.67 1.51
Ecuador -9.21 -5.90 -6.36
French Guiana -9.21 -9.99 -10.10
Guyana -9.21 -8.37 -8.57
Paraguay -9.21 -6.97 -7.43
Peru 11.24 7.06 8.53
Suriname -9.21 -8.32 -5.64
Uruguay -9.21 -10.50 -7.90
Venezuela -9.21 -11.01 -5.29
Table S.3: Numerical values of ground-truth and (a) predicted Area-Harvested of citrus fruit (b) predicted Yield of citrus fruit, in 26 European countries using livestock data from the same 26 countries - a supplement to Figure 2.
(a) Predicted Area-Harvested of Citrus Fruit
Country Ground Truth FATTNN TCN ($\mathcal{Y}\sim\mathcal{X}$)
Austria 10.01 10.74 13.54
Bulgaria 9.36 11.81 13.65
Croatia 9.18 10.03 0.33
Cyprus 8.39 12.07 11.58
Czechia 10.07 9.54 7.65
Denmark 8.55 7.98 1.90
Estonia 9.99 10.53 13.61
Finland 12.14 14.00 11.41
France 12.44 11.48 9.40
Germany 10.07 11.60 10.08
Greece 9.65 12.03 12.86
Hungary 9.14 11.37 8.96
Ireland 10.84 12.24 9.91
Italy 11.82 11.87 9.86
Latvia 10.71 10.94 5.96
Lithuania 9.75 8.53 4.28
Luxembourg 10.05 8.26 3.16
Malta 6.42 -1.43 -0.02
Netherlands 9.65 10.44 7.39
Poland 11.98 12.27 15.52
Portugal 12.61 12.11 22.19
Romania 12.02 14.08 18.88
Slovakia 9.00 6.94 -0.97
Slovenia 8.20 6.61 0.83
Spain 11.17 7.22 14.15
Sweden 10.10 7.71 13.22
(b) Predicted Yield of Citrus Fruit
@@
Figure S.4: Comparisons between ground-truth values and predicted pick-up and drop-off volumes using various methods. This plot is an enlarged version of Figure 3 of the main context. “District A to B” denotes the traffic volume that passengers were picked up in District A and dropped off in District B. The district numbering is assigned according to the Manhattan district map shown in Figure S.3. Left: Midtown Manhattan; Right: Downtown Manhattan.
Figure S.5: Comparisons between ground-truth values and predicted outputs for selected samples in FMRI dataset. From left to right: Ground truth, FATTNN, TCN, LSTM, LSTM, TRL, and Conv-TT-LSTM,.